FliC is a protein involved in bacterial flagella motion. It has a molecular weight of 43,545 Da and an extinction coefficient of 48,299. You have been able to purify this protein, and find that your 3 mL sample has an OD280 of 0.592. Length of the cuvette is 1.

What is the concentration of FliC?

A) 12.3 uM

How many nanomoles of FliC do you have?

A) 36.8 nM

Im not sure how they got 12.3 uM and then to convert that to nanomoles to get 36.8 nM.
Answer #1

Molecular Weight = 43545 Da: 43545 x 1.66x102 mg Protein concentration - OD 280 Extinction coeff x Cuvette length = 0.592 1.1

Molecular Weight = 43545 Da: 43545 x 1.66x102 mg Protein concentration - OD 280 Extinction coeff x Cuvette length = 0.592 1.1X105 48299X Number of moles of protein = Given mass Mol. mass = 1.0x10s 43545 81.66X1021 = 1.38*10 M = 1.38 x 10" MM No of molecules in IM = 6.023x10² then 1.38x104 MM contains = 6.023x10 x 1.384104 : 8.3 81027 23

1.38x104 uM = 1.38 x 104 x10-3 nM = 1.380 nM